The sum of two numbers is 400. One of them is 20 more than the other. Find the larger number.
Found 3 solutions by Alan3354, macston, stanbon:
Answer by Alan3354(69443)

The sum of two numbers is 400. One of them is 20 more than the other. Find the larger number.
=========================
x + y = 400
x - y = 20
------------------- Add

Answer by macston(5194)

Two numbers, one is 20 more than the other
X is the larger number, the other is X-20
The sum of the two numbers is 400
X+(X-20)=400
2X-20=400 Add 20 to each side
2X-20-+20=400+20
2X=420 divide both sides by 2
X=210 ANSWER the larger number is 210
CHECK:
the smaller number is 20 less
small=210-20=190
sum is 400: smaller+larger=400
190+210=400
400=400

Answer by stanbon(75887)

The sum of two numbers is 400. One of them is 20 more than the other. Find the larger number.
----------------
smaller:: x
larger::: x+20
------
Equation:
x + x+20 = 400
2x = 380
x = 190 (smaller #)
x+20 = 210 (larger #)
---------------
Cheers,
Stan H.
---------------